# RFNNS: Robust Fixed Neural Network Steganography with Popular Deep Generative Models

### v1

@@
Abstract.
Image steganography is a technique that conceals secret information in a cover image to achieve covert communication. Recent research has demonstrated that Fixed Neural Network Steganography (FNNS) exhibits significant practical advantages, as it enables stable and efficient steganographic embedding and extraction without requiring neural network training. However, the stego image generated by existing FNNS methods suffers from considerable distortion and exhibits poor robustness, severely reducing the security and practicality of steganography. To address the aforementioned issues, we propose a Robust Fixed Neural Network Steganography (RFNNS). In RFNNS, we introduce a texture-aware localization technique to add perturbations carrying secret image information to complex texture areas that are less perceptible to the human eye, thereby ensuring the quality of the stego image. To enhance robustness, a robust steganographic perturbation generation (RSPG) strategy is designed, which enables slight perturbations to be accurately decoded even after common image attacks. Subsequently, the generated robust perturbations are combined with the AI-generated cover image to produce the stego image. The receiver only needs to share the secret key and employ the same decoding network structure to accurately extract the secret image from the attacked stego image. Experimental results demonstrate that RFNNS achieves enhanced performance in terms of security, including imperceptibility and anti-steganalysis performance. Furthermore, RFNNS demonstrates superior robustness against common image attacks, such as JPEG compression, Gaussian noise, and contrast adjustment, across diverse embedding capacities, outperforming existing SOTA FNNS methods.
Robust steganography, Fixed neural network, Deep generative models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; XXX; XXXisbn: 978-1-4503-XXXX-X/2018/06submissionid: ccs: Security and privacy Security servicesccs: Security and privacy Human and societal aspects of security and privacy
1. Introduction
@@
With the rapid development of generative AI, the widespread application of generated content, such as images, has become increasingly prevalent in daily life, raising significant concerns about data security. Steganography (Mao et al., 2024; Li et al., 2024a; Wu et al., 2024; Kombrink et al., 2024; Steinebach, 2024; Meng et al., 2025), a critical information hiding technique (Liu et al., 2024; Wang et al., 2024; Zhu et al., 2024), ensures covert communication by embedding secret information in carriers such as images while remaining undetectable to humans and machine eavesdroppers, effectively safeguarding data security.
@@
Traditional image steganography relies on the Syndrome Trellis Codes (STCs) framework (Filler et al., 2011), which minimizes distortion to ensure security. Numerous distortion measurement functions (Holub and Fridrich, 2012, 2013; Li et al., 2014) have been designed on the basis of this framework. Recent advancements in deep neural networks (DNNs) have transformed steganography into a data-driven and learning-based approach (Baluja, 2017; Jing et al., 2021; Chen et al., 2022). However, this method faces two significant challenges: (1) it requires substantial data and computational resources to train effective neural networks; (2) the need to transmit trained models between senders and receivers prior to covert communication not only incurs storage overhead but also heightens the risk of detection by eavesdroppers, thereby compromising security.
Figure 1. The Process of Sending and Extracting in RFNNS.
@@
To avoid training and transmission of steganographic networks, researchers have employed Fixed Neural Networks (FNNs) (Ghamizi et al., 2021; Kishore et al., 2021; Luo et al., 2023; Li et al., 2024b) to embed and extract information. This approach leverages adversarial perturbations to modify the cover image such that the stego image can trigger a fixed-parameter decoding network to output the secret information. Covert communication can be achieved by sharing only the fixed decoding network architecture and the random seed to initialize the weights between the sender and the receiver. Nevertheless, existing FNNS methods are currently characterized by poor robustness against common image attacks, low stego image quality, and unsatisfactory anti-steganalysis performance. These limitations severely restrict the further development of this technology.
@@
In response to the aforementioned challenges, we propose an RFNNS method. Unlike previous FNNS methods (Ghamizi et al., 2021; Kishore et al., 2021; Luo et al., 2023; Li et al., 2024b), the perturbation embedded in our approach is not global but localized within selected regions. We propose a texture-aware localization technique that introduces perturbations carrying secret information into regions with high textural complexity that are less perceptible to the human eye. In addition, we devise a Robust Steganographic Perturbation Generation (RSPG) strategy that synthesizes perturbations resilient to a variety of common image attacks while keeping the distortion introduced into the stego images negligibly low. In practical applications, the receiver employs the shared secret key to access the meticulously designed decoding network we have developed, thereby reliably extracting the secret information. The sending and extraction process is shown in the Fig. 1.
An experiment was devised to thoroughly evaluate RFNNS relative to existing FNNS methods, including visual quality, robustness, and anti‑steganalysis performance. The results show that RFNNS outperforms the state-of-the-art in all three aspects. In terms of robustness, the secret images recovered by RFNNS under Gaussian‑noise attacks attain an average peak signal‑to‑noise ratio (PSNR) approximately 4 dB higher than that of competing methods. Across all contrast adjustment scenarios evaluated, it yields an average SSIM improvement of around 2-fold, demonstrating superior robustness. These gains stem primarily from the RSPG strategy, which safeguards the various embedded perturbations attacks. Because the texture‑aware localization technique confines perturbations to limited regions and narrows the gap between the stego and cover images, the proposed method consistently achieves a pronounced advantage in anti‑steganalysis performance, irrespective of whether the embedding rate is high or low. Due to its combined strengths in visual quality, robustness and anti‑steganalysis performance, RFNNS offers greater practical value for real‑world covert communication.
@@
  • We propose a texture-aware localization technique that embeds perturbations carrying secret information into high-texture-complexity regions, which are less perceptible to the human eye. This effectively reduces the distortion of the cover image caused by the perturbations.
@@
  • A robust steganographic perturbation generation (RSPG) strategy is designed to actively simulate potential attack scenarios that images may encounter during transmission. This strategy ensures that high-quality secret images can still be reliably extracted from the stego image after it has been subjected to common image attacks.
  • The Robust Fixed Neural Network Steganography (RFNNS) method is capable of accurately extracting a secret image through a meticulously designed fixed decoding network, even after the stego image has been subjected to common image attacks. Experiments demonstrate that, compared with other state-of-the-art FNNS methods, the RFNNS method exhibits superior performance in terms of visual quality, anti‑steganalysis performance, and robustness.
2. Related Work
2.1. Traditional Image Steganography
In traditional image steganography, hand-crafted embedding algorithms are typically employed to subtly modify the cover image, allowing secret information to be embedded while preserving visual quality to the greatest possible extent. Traditional image steganography methods can be broadly classified into spatial domain (Chan and Cheng, 2004; Pan et al., 2011) and transform domain (Westfeld, 2001) approaches. To further enhance the undetectability of stego images, researchers have proposed adaptive image steganography techniques (Holub and Fridrich, 2012). Adaptive steganography methods are often realized within a distortion coding framework, in order to minimize a defined distortion function resulting from data embedding. One of the most influential adaptive steganographic frameworks was presented in (Filler et al., 2011), using Syndrome-Trellis Codes (STCs) to encode secret information. Subsequently, various adaptive methods have been developed, each customized according to specific distortion functions (Holub and Fridrich, 2013; Li et al., 2014). Typically, to maintain a high level of undetectability, these methods restrict the embedding capacity to no more than 0.5 bits per pixel (bpp). In an effort to improve the robustness of the generated stego images for real-world applications, researchers have introduced robust steganography (Zeng et al., 2023; Tao et al., 2018; Cheng et al., 2025). Nevertheless, this approach remains constrained by critical challenges, notably its inherently limited embedding capacity and insufficient robustness against common yet potent image attacks, thereby diminishing its practical applicability.
@@ 2.1. Traditional Image Steganography
Figure 2. RFNNS framework: (a): Alice (The Sender) employs the proposed texture-aware localization technique to identify embedding regions corresponding to the perturbation. A robust steganographic perturbation generation (RSPG) strategy is then utilized to incorporate this perturbation into the AI-generated cover image, guided by a shared key, thereby producing the stego image. (b): The eavesdropping and potential image attacks that a stego image may encounter during transmission over a public channel. (c): Bob (The Receiver) first reconstructs the original cover image using the shared key to isolate the perturbation from the stego image. Subsequently, he decodes the secret image using the same decoding network. (d): Framework of Fixed Random Decoding Network.
2.2. DNN-based Image steganography
Recent advances in DNN-based (Szegedy et al., 2013; Yi et al., 2016; Cichy and Kaiser, 2019) steganography have enabled efficient embedding and extraction of secret information within cover images through specialized encoder and decoder networks. Zhu et al. (Zhu et al., 2018) introduced an end-to-end autoencoder-based image steganography method. SteganoGAN (Zhang et al., 2019) subsequently expanded this technology, achieving an embedding capacity of up to 6 bpp. Wei et al. (Wei et al., 2022) first proposed directly generating stego images using GANs (Goodfellow et al., 2014) without relying on separate cover images. Huang et al. (Huang et al., 2023) optimized GAN-based steganography by introducing an adaptive GAN parameter update strategy. Tang et al. (Tang et al., 2024) decoupled adversarial attack and information hiding processes, enhancing the extensibility of the method to embed information in various multimedia carriers. Peng et al. (Peng et al., 2024) utilized the inverse process of steganographic diffusion models to design a comprehensive encoding strategy that integrates the generation of steganographic space and the embedding of data, effectively improving steganography robustness.
However, these methods typically require extensive training data and considerable computational resources. Moreover, trained steganographic networks are often large in scale, necessitating covert transmission to senders or receivers who lack any steganographic tools. To simplify this process, FNNS emerged, enabling secret data embedding and extraction by adding adversarial perturbations to cover images without additional training. Ghamizi et al. (Ghamizi et al., 2021) proposed a steganographic approach that takes advantage of a multi‑label-targeted evasion attack, encoding the secret payload as the class labels assigned to the adversarial images generated during the evasion process. Kishore et al. (Kishore et al., 2021) proposed a method that increases embedding capacity by expanding the output dimension of the decoding network and utilizes information loss to generate encrypted images. Luo et al. (Luo et al., 2023) introduced a key-based method to synchronize steganographic tools and prevent unauthorized data recovery. Li et al. (Li et al., 2024b) proposed a Cover-separable Fixed Neural Network Steganography (Cs-FNNS) method that combines secret information with adversarial perturbations and optimizes the perturbations using a steganographic perturbation search algorithm. However, FNN-based steganographic methods often exhibit poor robustness against common image attacks, which hinders the accurate recovery of hidden information. Furthermore, these methods typically produce stego images that suffer from poor visual quality and pronounced distortion, ultimately limiting their practical applicability.
2.3. Deep Generative Models
@@
In recent years, deep generative models—such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), Variational Autoencoders (VAEs) (Kingma et al., 2013), Flow Models (Kingma and Dhariwal, 2018), and Diffusion Models (Ho et al., 2020) have advanced rapidly. Training in large-scale datasets approximates complex data distributions and has been widely used in Artificial Intelligence Generated Content (AIGC), achieving impressive results in computer vision (Ho et al., 2022), natural language processing (Brown et al., 2020), privacy protection (Yuan et al., 2022), and biological sciences (Zeng et al., 2022). For example, Rombach et al. (Rombach et al., 2022) employed diffusion models to synthesize high-resolution images; Ganev et al. (Ganev et al., 2024) integrated multiple models with differential privacy (DP) synthetic data generation techniques to mitigate privacy risks; and Lai et al. (Lai et al., 2025) applied deep generative models for drug molecule design, accelerating drug development and application. AIGC has also been utilized in information hiding. In RFNNS, by simply sharing a key and a prompt, both sender and the receiver can employ a generative model to produce identical cover images, thereby enabling precise localization of complex texture perturbations. Subsequently, on the basis of identified perturbation regions, we employ a RSPG strategy to embed the perturbations into specific areas of the generated cover image, thus producing a stego image. By integrating these techniques, our solution effectively enhances the practicality of covert communication.
@@
Figure 3. The texture-aware localization technique framework of the proposed method.
3. The Proposed Method
@@
3.1. Framework of the Proposed Scheme
In this study, we propose a novel steganography, called Robust Fixed Neural Network Steganography (RFNNS). For ease of description, the relevant symbols are shown in Table 1. Let $X_{c}$ represent an AI-generated RGB cover image, with $H_{c}$ and $W_{c}$ denoting its height and width, respectively. The secret image to be transmitted, denoted as $S$, is also an RGB image with height $H_{s}$ and width $W_{s}$. According to the framework depicted in Fig. 2, on the sender side, we input a secret key $k_{c}$ and a shared prompt into a pre-trained deep generative model $G(\cdot)$ to generate the cover image $X_{c}$.
(1) $X_{c}=G(k_{c},prompt)$
A texture-aware localization technique is employed to identify embedding regions within the cover image. Subsequently, the secret image is transformed into subtle perturbations denoted as $\delta$ using a robust steganographic perturbation generation (RSPG) strategy with a fixed decoding network. These perturbations are iteratively updated in response to various potential attacks. The refined robust perturbations are then embedded into predetermined regions of the cover image, ultimately generating the stego image.
(2) $\text{De}[k_{w}](\delta^{\prime})=S^{\prime}$
Table 1. Notations
@@ 3.1. Framework of the Proposed Scheme
$S’$ ’ Recoverd Secret Image $\in[0,1]^{H_{s}\times W_{s}\times 3}$
$G(\cdot)$ Deep Generative Model
@@
3.2. Texture-aware Localization
Existing FNNS methods typically encode secret information by uniformly embedding perturbations throughout the cover image, neglecting the substantial variations in texture complexity among different regions of the image. This uniform embedding strategy often leads to reduced visual quality and diminished anti-steganalysis performance. Considering human visual perception, subtle perturbations introduced into regions with high textural complexity are less likely to be perceptible. Therefore, embedding perturbations specifically within regions of higher textural complexity can effectively reduce the overall perturbation scale, thus enhancing the visual quality and anti-steganalysis performance.
As illustrated in Fig. 3, in practice, the cover image is initially partitioned into multiple equal-sized blocks of dimensions $b_{s}$×$b_{s}$. Subsequently, the texture complexity $O$ is computed for each block, and perturbations are introduced into the blocks whose complexity $O$ exceeds a predefined threshold $T$. We employ the Local Binary Pattern (LBP) (Ojala et al., 2002) method to quantify the $O$ of each block. For every pixel $p(i,j)$ in an image block, the corresponding LBP value is calculated by comparing the grayscale intensity of the central pixel with its eight neighboring pixels. The binary value $b_{k}$ for each neighbor pixel $p(i+dy,j+dx)$ is defined as follows:
(3) $b_{k}=\begin{cases}1,&p(i+dy,j+dx)\geq p(i,j),\\ 0,&p(i+dy,j+dx)<p(i,j)\end{cases}$
where $(dy,dx)$ represents the offset of each neighboring pixel relative to the central pixel, and $k$ $(k=0,1,\dots,7)$ denotes the neighbor index, arranged from left to right and then top to bottom. Following the LBP method described in (Pietikäinen, 2010), the resulting set of binary values $b_{k}$ is used to construct an LBP histogram $H(e)$. This histogram is subsequently normalized, yielding the probability distribution $P(e)$, from which we calculate the texture complexity $O$ as the entropy:
(4) $O=-\sum_{l=0}^{255}P(e)\log_{2}\left[P(e)+\epsilon\right]$

### v2

# RFNNS: Robust Fixed Neural Network Steganography with Universal Text-to-Image Models

@@
Abstract
With the rapid development of generative AI, image steganography has garnered widespread attention due to its unique concealment. Recent studies have demonstrated the practical advantages of Fixed Neural Network Steganography (FNNS), notably its ability to achieve stable information embedding and extraction without any additional network training. However, the stego images generated by FNNS still exhibit noticeable distortion and limited robustness. These drawbacks compromise the security of the embedded information and restrict the practical applicability of the method. To address these limitations, we propose Robust Fixed Neural Network Steganography (RFNNS). Specifically, a texture-aware localization technique selectively embeds perturbations carrying secret information into regions of complex textures, effectively preserving visual quality. Additionally, a robust steganographic perturbation generation (RSPG) strategy is designed to enhance the decoding accuracy, even under common and unknown attacks. These robust perturbations are combined with AI-generated cover images to produce stego images. Experimental results demonstrate that RFNNS significantly improves robustness compared to state-of-the-art FNNS methods, achieving an average increase in SSIM of 23% for recovered secret images under common attacks. Furthermore, the LPIPS value of recovered secrets images against previously unknown attacks achieved by RFNNS was reduced to 39% of the SOTA method, underscoring its practical value for covert communication. The code is available at https://github.com/edu-yinzhaoxia/RFNNS-Robust-Fixed-Neural-Network-Steganography-with-Universal-Text-to-Image-Models
1 Introduction
@@
With the rapid development of generative AI, the widespread application of generated content has become increasingly prevalent in daily life, raising significant concerns about data security. Steganography (lan2023robust; li2024purified; zhang2021universal; kombrink2024image; meng2025robust), a critical information hiding technique (yang2023autostegafont; xue2025physical; ji2025speech), ensures covert communication by embedding secret information in carriers such as images while remaining undetectable to humans and machine eavesdroppers, effectively safeguarding data security.
@@
Traditional steganography employs simple schemes such as least significant bit (LSB) replacement (van1994digital). Adaptive steganography selects suitable regions to modify during embedding. Recent advancements in deep neural networks (DNNs) have transformed steganography into a data-driven and learning-based approach (baluja2017hiding; jing2021hinet; chen2022hiding). However, this method faces two significant challenges: (1) it requires substantial data and computational resources to train effective neural networks; (2) the need to transmit trained models between senders and receivers prior to covert communication not only incurs storage overhead but also heightens the risk of detection by eavesdroppers, thereby compromising security.
Figure 1: The process of sending and extracting in RFNNS.
@@
To avoid training and transmission of steganographic networks, researchers have employed Fixed Neural Networks (FNNs) (ghamizi2021evasion; kishore2021fixed; luo2023securing; li2024cover) to embed and extract information. This approach leverages adversarial perturbations to modify the cover image such that the stego image can trigger a fixed-parameter decoding network to output the secret information. Covert communication can be achieved by sharing only the fixed decoding network architecture and the random seed to initialize the weights between the sender and the receiver. Nevertheless, existing FNNS methods are currently characterized by poor robustness against common image attacks, low stego image quality, and unsatisfactory anti-steganalysis performance. These limitations severely restrict the further development of this technology.
@@
In response to the aforementioned challenges, we propose an RFNNS method. Unlike previous FNNS methods (ghamizi2021evasion; kishore2021fixed; luo2023securing; li2024cover), the perturbation embedded in our approach is not global but localized within selected regions. We propose a texture-aware localization technique that introduces perturbations carrying secret information into regions with high textural complexity that are less perceptible to the human eye. In addition, we devise a Robust Steganographic Perturbation Generation (RSPG) strategy that synthesizes perturbations resilient to a variety of common image attacks while keeping the distortion introduced into the stego images imperceptibly low. In practical applications, the receiver employs the shared secret key to access the meticulously designed decoding network we have developed, thereby reliably extracting the secret information. The sending and extraction process is shown in the Fig. 1.
To evaluate the effectiveness of RFNNS in terms of visual quality, anti-steganalysis performance, and robustness, comprehensive benchmarking experiments were conducted against state-of-the-art FNNS methods. Experimental results indicate that RFNNS consistently achieves better performance compared with all baseline approaches. In particular, RFNNS demonstrates outstanding robustness generalization, maintaining high-quality recovery of secret images even under previously unknown attack scenarios. Experimental results demonstrate that RFNNS significantly improves robustness compared to state-of-the-art FNNS methods, achieving an average increase in SSIM of 23% for recovered secret images under common attacks. Moreover, under previously unknown attacks, the LPIPS value of recovered secrets achieved by RFNNS was reduced to 39% of the SOTA method, underscoring its significant robustness advantage.
@@
  • A texture‑aware localization technique is proposed to embed perturbations carrying secret information into regions of high texture complexity, which are less perceptible to the human eye. This effectively reduces the distortion of the cover image caused by the perturbations.
@@
  • A RSPG strategy is designed to actively simulate potential attack scenarios that images may encounter during transmission. This strategy ensures that high quality secret images can still be reliably recovered from the stego image even after it has been subjected to common or previously unknown image attacks.
  • Leveraging its meticulously designed fixed decoding network, RFNNS reliably recovers secret images even under common attacks. In addition, it surpasses leading FNNS baselines in visual quality, anti‑steganalysis performance, and robustness.
2 Related Work
2.1 Traditional Image Steganography
Traditional image steganography generally relies on manually designed algorithms to subtly embed secret information into cover images while maintaining their visual quality. Traditional image steganography methods can be broadly classified into spatial domain (chan2004hiding; pan2011image) and transform domain (westfeld2001f5) approaches. To further enhance the undetectability of stego images, researchers have proposed adaptive image steganography techniques (holub2012designing). Adaptive steganography operates within a distortion coding framework, epitomised by the Syndrome Trellis Codes (STCs) scheme of Filler et al. (filler2011minimizing) and later variants that fine‑tune the distortion metric for different covers (holub2013digital; li2014new). To remain hidden, these methods cap the payload at roughly 0.5 bpp. Robust steganography aims to resist channel degradations (zeng2023robust; tao2018towards; CHENG2025127598), yet it still struggles with limited capacity and vulnerability to routine image attacks.
@@ 2.1 Traditional Image Steganography
Figure 2: RFNNS framework: (a): Alice (The Sender) employs the proposed texture-aware localization technique to identify embedding regions corresponding to the perturbation. A RSPG strategy is then utilized to incorporate this perturbation into the AI-generated cover image, guided by a shared key, thereby producing the stego image. (b): The eavesdropping and potential image attacks that a stego image may encounter during transmission over a public channel. (c): Bob (The Receiver) first reconstructs the original cover image using the shared key to isolate the perturbation from the stego image. Subsequently, the same decoding network is employed to recover the secret image. (d): Framework of Fixed Random Decoding Network.
2.2 DNN-based Image steganography
Deep learning image steganography has moved from the pioneering end‑to‑end autoencoder of Zhu et al. (zhu2018hidden), through the SteganoGAN 6 bpp framework (zhang2019steganogan), to the recent StegFormer, which embeds multiple secrets in a single cover at up to 96 bpp while preserving high fidelity and robustness (ke2024stegformer).
However, these methods generally require extensive training data and computational resources, resulting in large network sizes challenging for covert transmission. FNNS emerged to simplify this process, embedding and extracting secret data through adversarial perturbations without additional training. Ghamizi et al. (ghamizi2021evasion) utilized multi-label evasion attacks for secret encoding. Kishore et al. (kishore2021fixed) increased payload by widening the decoder’s output channels and shaping perturbations via an information loss term. Luo et al. (luo2023securing) added a shared key to align sender and receiver, blocking unauthorized extraction. Li et al. (li2024cover) combined adversarial perturbations with steganographic search optimization. Nonetheless, FNNS techniques commonly face poor robustness against typical image attacks and significant visual distortions, which limits their practicality.
2.3 Universal Generative Text-to-Image Models
@@
In recent years, universal text-to-image models—such as Stable Diffusion XL(SDXL) (ICLR2024_081b0806), Stable Cascade Model (Pernias2024WrstchenAE), and Latent Diffusion Model (rombach2022high), have advanced rapidly. Training in large-scale datasets approximates complex data distributions and has been widely used in AIGC, achieving impressive results in computer vision (ho2022cascaded), natural language processing (brown2020language), privacy protection (tang2024once), and biological sciences (zeng2022deep; lai2025deep). AIGC has also been utilized in information hiding. RFNNS lets the sender and receiver regenerate an identical AI‑generated cover image from a shared key and prompt, then pinpoint high‑texture regions and through an RSPG strategy, embed localized perturbations, yielding a stego image that enhances practical covert communication.
@@
Figure 3: The texture-aware localization technique framework of the proposed method.
3 The Proposed Method
@@
3.1 Framework of the Proposed Scheme
In this study, we propose a novel steganography, called RFNNS. For ease of description, the relevant symbols are shown in Table 1. Let $X_{c}$ represent an AI-generated cover image, with $H_{c}$ and $W_{c}$ denoting its height and width, respectively. The secret image to be transmitted, denoted as $S$, is also an RGB image with height $H_{s}$ and width $W_{s}$. According to the framework depicted in Fig. 2, on the sender side, we input a secret key $k_{c}$ and a shared prompt into a pre-trained universal text-to-image model $G(\cdot)$ to generate the cover image $X_{c}$.
$X_{c}=G(k_{c},prompt)$ (1)
@@ 3.1 Framework of the Proposed Scheme
$\delta$’ Recoverd Micro Perturbation $\in[0,1]^{H_{\delta}\times W_{\delta}\times 3}$
@@
$S$’ Recoverd Secret Image $\in[0,1]^{H_{s}\times W_{s}\times 3}$
$G(\cdot)$ Universal Text-to-Image Model
@@
Table 1: Notations
A texture-aware localization technique is employed to identify embedding regions within the cover image. Subsequently, the secret image is transformed into subtle perturbations denoted as $\delta$ using a RSPG strategy with a fixed decoding network. These perturbations are iteratively updated in response to various potential attacks. The refined robust perturbations are then embedded into predetermined regions of the cover image, ultimately generating the stego image.
$\text{De}[k_{w}](\delta^{\prime})=S^{\prime}$ (2)
3.2 Texture-aware Localization
Existing FNNS methods typically encode secret information by uniformly embedding perturbations throughout the cover image, neglecting the substantial variations in texture complexity among different regions of the image. This uniform embedding strategy often leads to reduced visual quality and degraded overall performance. Embedding perturbations solely in highly textured regions, where human vision is least sensitive, minimizes overall distortion and thus improves visual quality and anti‑steganalysis performance.
As illustrated in Fig. 3, in practice, the cover image is initially partitioned into multiple equal-sized blocks of dimensions $b_{s}$ × $b_{s}$. Subsequently, the texture complexity $O$ is computed for each block, and perturbations are introduced into the blocks whose complexity $O$ exceeds a predefined threshold $T$. We employ the Local Binary Pattern (LBP) (ojala2002multiresolution) method to quantify the $O$ of each block (chosen for its computational simplicity and efficiency, and because it outperformed alternatives in our experiments). For every pixel $p(i,j)$ in an image block, the corresponding LBP value is calculated by comparing the grayscale intensity of the central pixel with its eight neighboring pixels. The binary value $b_{k}$ for each neighbor pixel $p(i+dy,j+dx)$ is defined as follows:
$b_{k}=\begin{cases}1,&p(i+dy,j+dx)\geq p(i,j),\\ 0,&p(i+dy,j+dx)<p(i,j)\end{cases}$ (3)
where $(dy,dx)$ represents the offset of each neighboring pixel relative to the central pixel, and $k$ $(k=0,1,\dots,7)$ denotes the neighbor index, arranged from left to right and then top to bottom. Following the LBP method described in (pietikainen2010local), the resulting set of binary values $b_{k}$ is used to construct an LBP histogram $H(e)$. This histogram is subsequently normalized, yielding the probability distribution $P(e)$, from which we calculate the texture complexity $O$ as the entropy:
$O=-\sum_{l=0}^{255}P(e)\log_{2}\left[P(e)+\epsilon\right]$ (4)